OpenStudy (anonymous):

Really need help solving this step by step. I know the answer is 2 but i dont know how to get the answer by doing it on paper 2√+6√2+3√−−−−−−√

OpenStudy (anonymous):

uber,what about me?

OpenStudy (anonymous):

\[\frac{\sqrt{2}+\sqrt{6}}{\sqrt{2+\sqrt{3}}}\]

OpenStudy (anonymous):

@Cali: There's enough uber to go around for everyone. :P

OpenStudy (anonymous):

@Courtney: Just to double check: "Square root of 2 plus square root of 6, all divided by square root of 2 plus square root of 3", right?

OpenStudy (anonymous):

\[\frac{\sqrt{2}+\sqrt{6}}{\sqrt{2+\sqrt{3}}}\]

OpenStudy (anonymous):

correct! ive been trying to figure it out but it's taking me FOREVER

OpenStudy (chaise):

You need to multiply by the conjugate. \[(\sqrt{2}+\sqrt{6}\div \sqrt{2}+\sqrt{6})\times(\sqrt{2}-\sqrt{6}\div \sqrt{2}-\sqrt{6})\] Solve this and you should get your answer

OpenStudy (anonymous):

chaise-would you do the \[\sqrt{6}\div \sqrt{2} \] first?

OpenStudy (chaise):

Multiply the numberators together, multiply the denominators together.

OpenStudy (anonymous):

would you mind showing me step by step because i honestly dont remember a lot about the rules of square roots and stuff

OpenStudy (chaise):

When you multiply by the conjugate you are not changing the answer. When you multiply by sqrt 2 - sqrt 6/sqrt - sqrt 6 (which is equal to 1, any number divded by itself is 1) you are not changing the answer. I dont know how to type to explain it. Do you want to join me here so I can explain? http://www.twiddla.com/583170

OpenStudy (anonymous):

Courtney: I was going to tell you that the conjugate is the opposite addition/subtraction of the denominator. So, if it says (y+x), the conjugate is (y-x). So, just do what chaise said, but be sure you know how to find conjugates for future reference.

OpenStudy (anonymous):

so you can still do the conjugate thing even though the denominator has a square root over the whole thing?

OpenStudy (anonymous):

uber??

OpenStudy (anonymous):

do you have the answers choices?

OpenStudy (anonymous):

(sqrt2+sqrt6)(sqrt(sqrt2-sqrt3))=sqrt2(sqrt(sqrt2+sqrt3))+sqrt6(sqrt(sqrt2-sqrt3))

OpenStudy (anonymous):

no i dont